OpenStudy (anonymous):

ok I need help solving this, if you can give me step by step on solving this equation it would be great :) 2n-2(2n+4)=8-(12-2n)

OpenStudy (johnweldon1993):

So we have \[\large 2n - 2(2n + 4) = 8 - (12 - 2n) \]

OpenStudy (johnweldon1993):

We first need to deal with distributing....distribute that 2 into those parenthesis...what do you get?

OpenStudy (anonymous):

you minus 2 from both sides right? @johnweldon1993

OpenStudy (anonymous):

or do you subtract 2n-2 and the multiply into the parenthesis?

OpenStudy (johnweldon1993):

Not quite...to distribute, you multiply that 2 to all terms inside the parenthesis... \[\large 2n - 4n - 8 = 8 - (12 - 2n)\] Make sense? because -2 times 2n = -4n and -2 times 4 = -8